Developing and Analyzing the Quadratic Formula
This formula, which is called the quadratic formula, can be used to find the roots for any quadratic equation of the form ax2 + bx + c = 0. The formula can be used to solve a problem such as the following: The house addresses of two neighbouring houses differ by two. Their product is 85 263. What is each address?
Purpose:
Use the quadratic formula to find the roots of quadratic equations.
Procedure:
A. Create a quadratic equation that can be used to determine the house addresses in the problem.
B. Use the quadratic formula to find the roots of the quadratic equation
C. Verify the roots in Step B by using a method presented in focus F
D. Explain how the roots of the quadratic equation can be used to solve the problem.
I need help with all of the procedure. I completely stuck

Answer by stanbon(75887)

The house addresses of two neighbouring houses differ by two. Their product is 85 263. What is each address?
----------
1st: x-1
2nd: x+1
-------------------
Equation:
(x-1)(x+1) = 85,263
------
x^2 - 1 = 85,263
--------
x^2 = 85,264
---------
x = 292
-----
x-1 = 291
x+1 = 293
--------------------
Cheers,
stan H.
===================